OpenStudy (anonymous):

Fluorine-17 is formed in a fusion reaction that also releases a neutron. If one of the reactants is an alpha particle, what is the other reactant?

OpenStudy (jfraser):

turn those names into proper nuclear symbols

OpenStudy (anonymous):

\[X_{z}^{a} + He _{_{2}}^{4}\rightarrow F _{9}^{17} + n _{0}^{1}\] this is the equation of the reaction you want to find\[X _{z}^{a}\] Conservation of mass says that \[a + 4 = 17 +1\] conservation of charge says \[x + 2 = 9\] So x the atomic number of the target X is ...? and a the atomic mass of the target X is ...?